accelerates uniformly from a velocity of 1 m s-1 to a velocity of 5 m s-1 in 0.5 minutes . Calculate the displacement of the transporter .


1
Expert's answer
2021-02-19T10:31:01-0500

By definition, the average acceleration is given as:


"a = \dfrac{v_2-v_1}{t}"

where "v_1,v_2" are initial and final velocities respectively, and "t" is the time interval the object experienced the acceleration. Substituting our values, obtain:


"a = \dfrac{5m/s-1m/s}{0.5s} = 8m/s^2"

The distance traveled under the constant acceleration is given as:


"d = v_1t + \dfrac{at^2}{2}\\
d = 1m/s\cdot 0.5s + \dfrac{8m/s^2\cdot (0.5s) ^2}{2} = 1.5m"

Answer. 1.5 m.